# Learning to Be A Doctor: Searching for Effective Medical Agent Architectures

### v1

@@ 4.1 Experimental Setup
Figure 3: Workflow evolution over iterations. We visualize the diagnostic workflow at three key stages: the initial state (Iteration 0), an intermediate state after partial refinement (Iteration 3), and the final evolved workflow (Iteration 6). The evolution process demonstrates how the system incrementally improves its structure by adding conditional branches, parallel reasoning paths, and error-specific loops. These structural refinements enable more accurate and robust diagnostic reasoning through iterative self-adaptation.
Datasets. To evaluate the effectiveness of our proposed framework, we select the task of skin disease diagnostics as the primary testbed. This domain is particularly suitable for evaluating multi-agent architectures due to its inherent complexity: accurate diagnosis often requires interdisciplinary knowledge and a multi-step reasoning process involving visual inspection of skin lesions, patient history, and symptom-diagnosis correlations. We utilized two dermatological datasets to evaluate the workflow evolution:
SKIN Concepts Dataset. SKINCONDaneshjou et al. (2023) builds upon the Fitzpatrick 17k skin disease datasetGroh et al. (2022, 2021), including 3230 images annotated with 48 clinically relevant concepts, such as "plaque," "scale," and "erosion." These annotations span 113 types of skin diseases and are based on established clinical vocabulary used by dermatologists, ensuring their applicability across diverse disease processes.
Augmented Skin Conditions Image Dataset. Skin Conditions Naqvi (2023)contains a total of 2,394 augmented images across six different dermatological conditions, with each category comprising 399 images. This balanced structure makes it ideal for training machine learning models, particularly in the field of medical image analysis.
@@
Data Processing. During preliminary data preparation, we identified a subset of images within the datasets that depicted skin conditions located in private or sensitive body areas. Due to content moderation constraints, these images were rejected by the LLM and could not be processed during workflow execution. Additionally, we encountered another issue where approximately 20 images had broken URL links, preventing them from being downloaded. To maintain the integrity of our experiments and ensure consistent, reliable evaluation across all agent architectures, we excluded these samples from the dataset.
Datasets Split: SKIN Concepts Dataset: The training and validation sets contain 50 randomly sampled images each, and the test set includes 224 images, ensuring that all disease types are represented with two images per class. SKIN Conditions Dataset: Training and validation sets each have 50 images; the test set contains 120 images, with 20 per condition.
Implementation Details. We implemented an experimental process that incorporates training, validation, and testing phases.
After each training batch, the evolutionary mechanism adjusts the workflow structure. The model’s performance is then validated on a separate set. This cycle of training and evolution repeats until training completion, followed by a final evaluation of the test set.
Our framework is built on LangGraphInc. (2023). LangGraph provides us with some pre-packaged content, and we only need to develop our node designs on top of LangGraph. Additionally, we provide a tool list where you can add your custom tools. In the tool list, we have defined an image search tool for finding similar images. Image search tool utilizes the pre-trained CLIP(ViT-L/14)Radford et al. (2021) model to extract feature vectors without any further fine-tuning from images for data that is not part of the training, validation, and testing sets, storing these vectors in the PineconePinecone Systems (2023) database. When searching for similar images using the image tool search, it can directly match the feature vectors stored in this database.
Figure 4: Top-k diagnostic accuracy over workflow evolution iterations. This figure shows how the Top-k accuracy improves as the workflow undergoes iterative optimization. Each point represents the diagnostic performance after one round of workflow evolution. The increasing trend demonstrates that structural refinements and prompt updates effectively enhance diagnostic capability. The curve eventually converges, indicating that the system stabilizes once a sufficiently optimized workflow is discovered. The results in the figure were obtained using the GPT-4o.
Baselines. To establish a comparative framework for evaluating the effectiveness of our workflow evolution methodology, we implemented two baseline approaches for image-based dermatological diagnosis. The first baseline utilizes GPT-4o as a single-agent model to perform direct diagnosis of input images, without any intermediate reasoning or workflow adjustments. This provides a straightforward benchmark for assessing diagnostic accuracy.
The second baseline augments GPT-4o with a Chain-of-Thought (CoT) prompting mechanism. This approach guides GPT-4o to reason through intermediate steps before producing the final diagnosis. Although CoT enhances interpretability compared to direct diagnosis, it does not leverage adaptive or evolving workflows, which are central to our proposed methodology.
Figure 5: Per-disease consensus accuracy (cons@64) across different methods.This figure illustrates the prediction stability of each method under six different disease categories, evaluated via majority-vote consensus over 64 samples per class. Our method consistently achieves higher cons@64 scores, indicating improved diagnostic robustness and intra-class agreement across cases.
The third baseline involves a "round table" discussion setup, which is a type of multi-agent system configuration. In this approach, multiple agents collaborate and deliberate around a common task or theme to achieve a more efficient or accurate solution. Each agent may employ distinct models or algorithms and share information with one another, akin to a round table meeting, to integrate diverse insights and optimize the final outcome.
These baselines allow for quantifying the benefits of workflow evolution in terms of generalization ability and diagnostic performance.
LLM Method Skin Concepts Accuracy (%) Skin Conditions Accuracy (%) Top-1 Top-3 Top-5 Top-1 Top-3 Top-5 GPT-4o IO 20.27 30.63 36.04 50.83 78.33 86.67 CoT Wei et al. (2022) 18.47 28.83 33.78 55.83 76.67 82.50 Round Table Chen et al. (2024b) 21.17 27.93 32.43 45.83 75.83 80.83 Ours 29.28 40.09 50.45 90.83 95.00 100.00 GPT-4o-mini IO 11.71 20.72 23.87 27.50 69.17 80.83 CoT Wei et al. (2022) 6.31 15.32 24.32 22.50 65.00 84.17 Round Table Chen et al. (2024b) 10.81 19.82 23.42 25.83 70.00 78.33 Ours 13.51 21.62 24.77 45.83 74.17 85.83
@@
Table 1: Top-k diagnostic accuracy (%) of different methods using GPT-4o and GPT-4o-mini on skin concepts and skin conditions. Higher values indicate better performance.
@@
Table 2: Ablation results of different operations. Each row shows the Top-1, Top-3, and Top-5 accuracies when a specific operation is disabled. Values in parentheses indicate accuracy drop compared to the full model Top-1(29.28.%), Top-3(40.09%), and Top-5(50.45%).
Evaluation Metrics. To evaluate the diagnostic models, we adopted accuracy-based metrics across Top-1, Top-3, and Top-5 predictions. Each diagnostic attempt produces a ranked set of predictions based on the model’s confidence scores, and the correctness of these predictions is compared against ground truth labels.
During the validation phase, the Top-1, Top-3, and Top-5 accuracies are measured after each evolution iteration, enabling us to observe how these metrics improve as the workflow evolves. In the test phase, the final accuracies obtained with our workflow evolution method are compared against those of the baselines, demonstrating the advantages of dynamic and adaptive workflows.
Following Wang et al. (2023), we adopt the majority voting consensus metric, denoted as cons@64, which reflects how often a method arrives at a consistent diagnosis across multiple inputs of the same class (n=64). This metric offers a proxy for model robustness and reliability under input variance and sampling noise.
4.2 Analysis of Evolutionary Processes
Evolution Visualization. If the workflow nodes are directly input to the LLM in text form, the LLM may easily overlook structural relationships such as loops and conditional branches. To better understand how our workflow evolves over time, we visualize the structure at three representative stages: the initial workflow (Iteration 0), an intermediate state (Iteration 3), and the final evolved workflow (Iteration 6). These visualizations reveal how the system progressively incorporates structural refinements, such as conditional branches, parallel reasoning paths, and error-specific loops. Figure 3 presents a comparative view of these stages.
Quantitative Analysis To quantitatively evaluate the effectiveness of our workflow evolution framework, we track the diagnostic performance—measured by Top-1 accuracy, Top-3 accuracy and Top-5 accuracy—across multiple evolution iterations. At each iteration, the system refines its workflow based on prior diagnostic errors and structural improvements, allowing it to incrementally enhance its reasoning capabilities.
Figure 4 illustrates how the Top-1 accuracy, Top-3 accuracy, and Top-5 accuracy change as the workflow evolves. The straight line represents the accuracy of directly inputting images into the LLM for diagnosis. It can be observed that the accuracy of the evolved workflow is always higher than the IO line. We observe a consistent upward trend during the early iterations, indicating that the refinement of workflow structure and prompt configurations leads to improved diagnostic performance. The accuracy eventually converges, suggesting that the system has reached a stable and effective workflow configuration. This pattern demonstrates the framework’s ability to learn from experience and self-optimize over time.
Additionally, in the Skin Concepts dataset, both the training and validation sets consist of 50 randomly selected images, while the testing set includes images from all categories. Our workflow iteratively evolves on the training set and is validated on the validation set. Ultimately, it achieves good results on the testing set, which includes all categories. This demonstrates the generalization capability of our evolved workflow.
4.3 Performance Comparison Results
Diagnostic Accuracy Comparison. To comprehensively evaluate the effectiveness of our method in dermatological diagnostic tasks, we compare various prompting strategies across multiple state-of-the-art large language models (LLMs). Specifically, we report the Top-k diagnostic accuracy on both skin concepts and skin conditions using GPT-4o and GPT-4o-mini. As shown in Table 1, our approach consistently outperforms existing baselines, including IO, CoT Wei et al. (2022), and Round TableChen et al. (2024b), achieving superior results across all LLMs and evaluation metrics. As shown in Figure 1, our method achieves substantial gains in Top-k accuracy, highlighting its superior diagnostic capability and robustness across different models.
Workflow Stability. In addition to diagnostic accuracy and structural adaptability, we assess the stability of our evolved workflows by evaluating their consensus behavior across multiple independent samples. We report per-class consensus accuracy (cons@64) for each method using the GPT-4o model on the Skin Concepts dataset, as shown in Figure 5. Our method consistently achieves higher consensus scores across a wide range of conditions, including both common (e.g., eczema, psoriasis) and rare or visually complex diseases (e.g., impetigo, rosacea). This demonstrates not only the general robustness but also the fine-grained reliability of the evolved workflows.
4.4 Component Ablation Analysis
We conduct a comprehensive ablation study to investigate the role of different workflow modification actions in our evolution framework. Specifically, we categorize all actions into three main types: Add, Modify, and Remove. Each of these categories encompasses a set of fine-grained operations applied to either nodes or structural components (e.g., conditional, loop, parallel).
In this study, we conducted ablation experiments to evaluate the impact of different operations on the effectiveness of the workflow. Specifically, we selected three key operations from the set of possible actions: adding a tool node, modifying a node prompt, and removing a node. Firstly, the operation of adding a tool node is intended to acquire information about similar images. To assess the necessity of this operation, we removed it from the search space and observed the performance changes. Similarly, we applied the same treatment to the operations of modifying a node prompt and removing a node, excluding them from the search space to analyze each operation’s contribution to the overall system performance and effectiveness. These experiments will aid in understanding the role and significance of each operation in workflow optimization.
To assess their individual contribution, we disable one specific action at a time and examine the resulting Top-1 diagnostic accuracy on the Skin Concepts dataset. The results are summarized in Table 2, which reveals the relative importance of each operation in enabling effective performance improvements. On the one hand, notably, actions related to adding tool node and prompt modification exhibit a particularly strong impact on performance, underscoring their significance in workflow adaptability and optimization. On the other hand, we can see that removing nodes does not significantly impact accuracy. However, in practice, having too many nodes with similar functions can lead to decreased execution efficiency of the workflow.
@@ 5 Conclusion
Current dermatology-focused multi-agent systems are hindered by rigid, manually designed workflow frameworks that lack the adaptability to meet evolving diagnostic demands and diverse requirements. To overcome this limitation, we propose a node-level evolution framework that introduces flexibility through autonomous exploration and development of feasible workflows. By enabling dynamic evolution at the node level, our framework allows workflows to adapt organically, facilitating efficient responses to changing conditions and complex diagnostic scenarios. This approach significantly enhances flexibility by supporting modular updates, structural adjustments, and parallel processing while autonomously optimizing workflow paths. As a result, our framework offers a scalable and adaptive solution for improving diagnostic accuracy, operational efficiency, and responsiveness in dermatological multi-agent systems.

### v2

@@ 4.1 Experimental Setup
Datasets. To evaluate the effectiveness of our proposed framework, we conduct experiments on two dermatological image classification datasets Daneshjou et al. (2023); Naqvi (2023) to assess workflow evolution. Dermatology presents a challenging and representative domain for evaluating multi-agent architectures, as accurate diagnosis typically demands both specialized visual analysis and integrative reasoning. Below is a detailed description of the two datasets.
Skin Concepts Daneshjou et al. (2023) extends the Fitzpatrick 17k skin disease dataset Groh et al. (2022, 2021) and includes 3,230 images annotated with 48 clinically relevant skin concepts. For evaluation, we adopt a split of 50 images for training, 50 for validation, and 224 for testing, ensuring two test images per disease type to maintain balanced coverage across all categories.
Augmented Skin Conditions Naqvi (2023) consists of 2,394 synthetically augmented images across six dermatological conditions, with 399 images per class. For evaluation, we adopt a similar split of 50 training, 50 validation, and 120 test images, with 20 images per disease type, to ensure sufficient and balanced evaluation of diagnostic performance.
@@
During data processing, we observed that the LLM rejected a small number of images depicting sensitive body areas due to content moderation policies. Additionally, some image data with broken URLs could not be downloaded. To ensure experimental integrity and enable fair and consistent evaluation across all agent architectures, we excluded these samples from our experiments. In total, 100 images from the Skin Concepts Daneshjou et al. (2023) dataset were removed.
Figure 4: Top-k accuracy improves with workflow iterations, and convergence signals system stabilization using GPT-4o.
Baselines. To establish a comparative framework for evaluating the effectiveness of our workflow evolution methodology, we implemented three baseline approaches. The first baseline utilizes LLM to give a direct diagnosis of input images, without any intermediate reasoning or workflow adjustments. This provides a straightforward benchmark for assessing diagnostic accuracy. The second baseline augments the LLM with a CoTWei et al. (2022) prompting mechanism. This approach guides LLM to reason through intermediate steps before producing the final diagnosis. The third baseline adopts the Round Table Chen et al. (2024c) framework, a multi-agent collaboration paradigm inspired by human expert panels. In this setup, multiple agents take turns to refine a shared solution through iterative deliberation. Each agent reviews the reasoning history and outputs of previous agents, then contributes an updated response based on its own perspective. This process continues for a fixed number of iterations or until convergence, aiming to achieve a more comprehensive and accurate final decision.
Implementation Details. Our experimental setup includes training, validation, and testing phases. During training, the workflow evolves through iterative refinement based on diagnostic outcomes. Each updated workflow is validated after every batch to assess performance improvements. Once training completes, the final workflow is evaluated on the test set to measure overall performance.
Our framework extends LangGraph Inc. (2023) with customized node designs to support flexible and modular workflow construction. To facilitate extensibility, we introduce a configurable tool list that allows users to easily add or integrate their own custom tools. As a built-in example, we include an image search tool designed to find visually similar images. We use the pre-trained CLIP (ViT-L/14) Radford et al. (2021) model to extract feature vectors from images that are not included in the training, validation, or test sets. These vectors are then stored in the Pinecone Pinecone Systems (2023) vector database to support image similarity searches. When a new image is input, its feature vector is computed and compared against the stored vectors to enable fast and efficient similarity searches.
To evaluate the generality and robustness of our framework across diverse LLMs, we conduct experiments using the following models: GPT-4o-2024-05-13 Achiam et al. (2023), Claude-3-5-sonnet-20240620 Anthropic. (2024), and GPT-4o-mini-2024-07-18 Menick et al. (2024). The temperature parameter is set to 1.0 to encourage diverse reasoning paths, and the random seed is 42.
@@
Table 1: Top-k diagnostic accuracy (%) of different methods using GPT-4o, GPT-4o-mini and Claude 3.5 Sonnet on Skin Concepts and Skin Conditions.
LLM Method Skin Concepts Accuracy (%) Skin Conditions Accuracy (%) Top-1 Top-3 Top-5 Top-1 Top-3 Top-5 GPT-4o IO 20.27 30.63 36.04 50.83 78.33 86.67 CoT Wei et al. (2022) 18.47 28.83 33.78 55.83 76.67 82.50 Round Table Chen et al. (2024c) 21.17 27.93 32.43 45.83 75.83 80.83 Ours 29.28 40.09 50.45 90.83 95.00 100.00 GPT-4o-mini IO 11.71 20.72 23.87 27.50 69.17 80.83 CoT Wei et al. (2022) 6.31 15.32 24.32 22.50 65.00 84.17 Round Table Chen et al. (2024c) 10.81 19.82 23.42 25.83 70.00 78.33 Ours 13.51 21.62 24.77 45.83 74.17 85.83 Claude 3.5 Sonnet IO 17.12 24.77 26.13 40.00 68.33 75.83 CoT Wei et al. (2022) 14.86 22.07 24.32 36.67 63.33 73.33 Round Table Chen et al. (2024c) 15.77 22.52 26.58 35.00 66.67 74.17 Ours 28.83 35.14 38.29 95.83 98.33 99.17
Evaluation Metrics. To evaluate diagnostic models, we used Top-1, Top-3, and Top-5 accuracy metrics, comparing predictions against ground truth labels. During validation, these accuracies were tracked across workflow evolution to observe improvements, while in testing, final accuracies were compared with baselines to highlight the benefits of adaptive workflows.
In addition to accuracy, we assess the consistency of predictions across repeated inputs. We adopt the majority voting consensus metric, denoted as cons@64 Guo et al. (2025), which reflects how often a method arrives at a consistent diagnosis across multiple inputs of the same class (n=64). This metric offers a proxy for model robustness and reliability under input variance and sampling noise.
4.2 Results
Evolution Visualization. To better understand how our workflow evolves, we visualize the structure at three representative stages: the initial workflow (Iteration 0), an intermediate state (Iteration 3), and the final evolved workflow (Iteration 6). Figure 3 presents a comparative view of these stages. These visualizations reveal how the system progressively incorporates structural refinements, such as conditional branches, parallel reasoning paths, and error-specific loops.
Cost and Complexity Analysis. The results in Table 2 show that the framework evolves into more complex workflows by the end of training. The final workflows used in testing contain more nodes and branches than the training-phase averages, indicating an increase in both structural complexity and token consumption driven by the evolutionary process.
Table 2: Resource Consumption and Workflow Complexity. This table summarizes resource consumption and workflow complexity across different metrics during training (Evo) and testing phases.
Metric Train (Evo) Test
Number of Nodes 8.7 11
Number of Branches 4.0 5
Time (s) 210 621
Total Tokens (k) 361 349
Performance Over Iterations. To quantitatively evaluate the effectiveness of our workflow evolution framework, we track the diagnostic performance, measured by Top-1 accuracy, Top-3 accuracy, and Top-5 accuracy across multiple evolution iterations. At each iteration, the system refines its workflow based on prior diagnostic errors and structural improvements, allowing it to incrementally enhance its reasoning capabilities.
Figure 4 illustrates the iterative enhancement in accuracy as the workflow progresses, compared to the baseline accuracy achieved by directly inputting images into the LLM for diagnosis (IO line). The evolved workflow consistently surpasses the baseline performance, exhibiting an initial upward trend attributed to structural refinement and prompt optimization. This is followed by convergence as the system stabilizes into an effective configuration. This pattern underscores the framework’s capacity to learn from experience and self-optimize over time. In the context of the Skin Concepts dataset Daneshjou et al. (2023), the evolved workflow demonstrates robust generalization capabilities, achieving commendable results across all categories in the testing set.
Evolution Results Analysis. To comprehensively evaluate the effectiveness of our method in diagnostic tasks, we compare it against various baselines as shown in Table 1. Specifically, we report the Top-k diagnostic accuracy on dermatological tasks using the Skin Concepts Daneshjou et al. (2023) and Skin Conditions Naqvi (2023) datasets across three different LLMs.
Our approach consistently outperforms existing baselines, including IO, CoT Wei et al. (2022), and Round Table Chen et al. (2024c), across all LLMs and evaluation metrics. Our method achieves superior results in Top-1, Top-3, and Top-5 accuracies for both datasets. These results highlight the adaptability and robustness of our method across different LLMs and diverse medical diagnostic scenarios. The consistent superiority in Top-k accuracies underscores the effectiveness and reliability of our evolved workflow, making it a promising solution for improving diagnostic performance.
Figure 5: Per-disease consensus accuracy (cons@64) across different methods. This figure shows prediction stability across six disease categories, evaluated via majority-vote consensus over 64 samples per class. Our method achieves higher cons@64 scores, indicating greater diagnostic robustness and consistency.
Workflow Stability. In addition to diagnostic accuracy, we assess the stability of our evolved workflows by evaluating their consensus behavior across multiple independent samples. We report per-class consensus accuracy (cons@64 Guo et al. (2025)) for each method using the GPT-4o model on the Skin Concepts dataset Daneshjou et al. (2023), as shown in Figure 5. Our method consistently achieves higher consensus scores across a wide range of conditions.
4.3 Ablation Analysis
We conduct a comprehensive ablation study to evaluate the contribution of individual workflow modification actions in our evolution framework. Actions are classified into three categories: Add, Modify, and Remove. In the study, we assess the impact of three key operations: adding tool nodes, modifying node prompts, and removing nodes. By excluding each operation from the search space, we observe its contribution to overall performance. Table 3 shows that adding tool nodes and modifying prompts notably improve accuracy, while removing nodes has minimal impact on accuracy but enhances execution efficiency by reducing redundancy.
Table 3: Ablation results showing Top-k accuracies when specific operations are disabled. Parentheses indicate accuracy drops relative to the full model.
@@ 4.3 Ablation Analysis
We further evaluate the framework under three different integration levels: node-level (N), node + structural-level (N+S), and full-level (N+S+F). The results show that incorporating structural-level modifications leads to performance improvement, with the full-level framework achieving the best results. This confirms the effectiveness of hierarchical integration in supporting comprehensive workflow evolution.
Table 4: Ablation results for integrating different operation levels (N, N+S, N+S+F). The fully-leveled framework delivers the best performance.
Setting Top-1 Acc. (%) Top-3 Acc. (%) Top-5 Acc. (%) N 26.13 36.94 45.05 N+S 27.03 38.74 46.40 N+S+F 29.28 40.09 50.45
Together, these experiments provide a comprehensive understanding of how individual operations and integration levels contribute to performance, offering interpretable evidence for the design in our framework.
@@ 5 Conclusion
Current medical diagnostic systems are often constrained by rigid, manually designed workflows that lack the flexibility to adapt to evolving clinical demands and diverse diagnostic requirements. To address this limitation, we propose a workflow evolution framework that enables the autonomous exploration and iterative refinement of agent-based diagnostic workflows. To enable a more responsive and robust diagnostic workflow, our framework introduces dynamic adaptability at both the structural and prompt levels, allowing it to evolve in response to diagnostic errors. In summary, our framework provides a scalable and adaptive solution for advancing multi-agent systems in diagnostic tasks. It not only bridges the gap between static, expert-defined architectures and dynamic, self-improving methods, but also lays the groundwork for future adaptive diagnostic systems.